# Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation

### v2

@@ Abstract
To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG). While prior work focuses on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations. We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents. Theanine discards memory removal and manages large-scale memories by linking them based on their temporal and cause-effect relation. Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts in measuring memory-augmented dialogue agents. A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.    Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University
@@ 1 Introduction
Autonomous agents based on large language models (LLMs) have made great progress in various domains, including response generation (Zhou et al., 2023; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly. To facilitate such lifelong interaction, studies have proposed to improve dialogue agents’ ability to memorize and correctly recall past information (e.g., discussed topics) in long-term, multi-session conversations.
@@
A representative method is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023). However, the growing span of memories can hinder retrieval quality as conversations accumulate. Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), this common practice may cause severe information loss. As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update and causes improper RG. While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024). this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024). These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to effectively store large-scale past interactions without removing old memories? (ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
@@
Motivated by these, we propose to address the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span. We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework for facilitating lifelong dialogue agents.
@@
Starting from memory construction (Phase I), instead of stacking raw memory sentences as-is (Xu et al., 2022a), which may affect memory retrieval and also response quality due to the unstructured format of information (Mousavi et al., 2023; Chen et al., 2023), Theanine stores memories in a directed graph. In this graph, inspired by how humans naturally link new memories to existing ones of relevant events based on their relation (Bartlett, 1995), memories are linked using their temporal and cause-effect commonsense relations (Hwang et al., 2021). Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023). Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation (Madaan et al., 2024), such that they provide tailored information (Chae et al., 2023) for generating responses (Phase III). Our contributions are two-fold:
@@
  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline-augmented response generation for long-term conversations. Theanine outperforms baselines across automatic, LLM-based, and human evaluations of RG. Also, we confirm that Theanine leads to better retrieval quality, and its intermediate procedures align with human preference. To our knowledge, we are the first to model the concept of timelines in memory management and RG.
@@ 2 Methodologies
We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents, inspired by how humans store and retrieve memories for conversations (Figure 2):
@@ 2.2 Timeline Retrieval and Timeline Refinement (Phase II)
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] Based on our experiments, we empirically set $n$ to 1. Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$. In practice, $M_{re}$ are processed together, although processing them 1-by-1 yields the same result. Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
@@ 3 Experimental Setups
Table 1: Automatic evaluation of response quality (average of sessions). Session-specific results are in Appendix F.
@@ 3.1 Datasets of Long-term Conversations
There are few datasets of long-term, multi-session conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean. Recently, Jang et al. (2023) release a new dataset, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^8] https://www.xinli001.com/ which addresses long-term conversations in Mandarin regarding clinical scenarios.
We use MSC and CC for evaluation. We focus on English conversations, so we do not adopt DuleMon, CareCall, and Psychological QA and leave multilingual and clinical settings to future work.
@@ 3.2 Baselines
To evaluate Theanine, besides naive baselines using all past dialogues or memories, we include:
@@
Memory Update. We use LLMs to run a widely used updating algorithm proposed by Bae et al. (2022) at the end of each session. The algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
RSum-LLM. An LLM-only generative method, which recursively summarizes and updates the memory pool and accordingly generates responses (w/o a retrieval module) (Wang et al., 2023).
MemoChat. Proposed by Lu et al. (2023), it uses LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
COMEDY. Proposed by Chen et al. (2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation, and selects them to augment RG.
@@ 3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 4), (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity. In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3. For “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost the same in sessions 1-2 (no memory to update).
@@ 4 Evaluation Scheme 1: Automatic and Human Evaluations
To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation. We now present the main findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
@@
Table 1 reports agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020). Theanine, in both datasets, achieves superior response quality than diverse baselines. Despite slightly lower scores compared to Memory Retrieval in overlap-based metrics (i.e., B-4 and R-L) in MSC, Theanine largely outperforms it in terms of embedding-based metrics. Interestingly, including ours, methods w/o memory update generally yield higher scores, justifying our proposal towards an update-free memory management for lifelong dialogue agents.
@@
To gain deeper insights into our design, we show Theanine’s performance when relation-awareness in linking (Phase I-2) and Timeline Refinement (Phase II-2) are ablated. Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
In Table 2, we observe a ranking regarding contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement. We thus confirm the efficacy of constructing a memory graph with causal relations. Moreover, utilizing such a structure to collect timelines of relevant events yields better RG than conventional retrieval, despite the smaller $k$ (3 vs. 6) in initial retrieval. Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG. We leave it to future work.
(Finding 4) Humans and G-Eval both reveal that Theanine retrieves memories that are more helpful for response generation.
Beyond agent responses, we further investigate how different memory construction methods affect retrieved memories’ helpfulness in augmenting RG. Given the same current dialogues as queries for retrieval, Figure 5 (left; G-Eval; right: Human Eval) shows head-to-head comparisons (ours vs baselines) regarding whose retrieved memories benefit response generation more. We observe higher win rates of Theanine in all comparisons, especially in human evaluations, suggesting that ours can facilitate a more helpful memory-augmentation for RG.
@@
Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories can help us facilitate reliable lifelong dialogue agents.
For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting. In Figure 6, we find Theanine not only leading to a small number of contradictory responses (4%) but also showing the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines. We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses that more directly align with the past, which is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020). Also, such entail, non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
@@
As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, explaining its contribution to performance. Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG. Examples of Theanine’s phases and RG are in Appendix G.

### v3

# Towards Lifelong Dialogue Agents via Timeline-based Memory Management

@@ Abstract
To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG). While prior studies focus on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations. We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents. Theanine discards memory removal and manages large-scale memories by linking them based on their temporal and cause-effect relation. Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts when assessing agent performance in integrating past memories into RG. A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
Towards Lifelong Dialogue Agents via Timeline-based
Memory Management
Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.    Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University {ktio89, namyoung.kim, jinyeo}@yonsei.ac.kr
@@ 1 Introduction
Autonomous agents based on large language models (LLMs) have made significant progress in various domains, including response generation (Chae et al., 2024; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly. To facilitate such lifelong interaction, studies have proposed enhancing dialogue agents’ ability to memorize and accurately recall past information (e.g., discussed topics) in long-term, multi-session conversations.
@@
A representative approach is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023). However, the growing span of memories can hinder retrieval quality as conversations accumulate. Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), such common practice may cause severe information loss. As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update, resulting in improper RG. While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024). this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024). These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to store large-scale past interactions effectively without removing old memories? (ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
@@
Motivated by these, we propose addressing the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span. We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework for facilitating lifelong dialogue agents.
@@
Starting from memory construction (Phase I), instead of stacking raw memory sentences as-is (Xu et al., 2022a), which may affect memory retrieval and also response quality due to the unstructured format of information (Mousavi et al., 2023; Chen et al., 2023), Theanine stores memories in a directed graph. In this graph, inspired by how humans naturally link new memories to existing ones of relevant events based on their relation (Bartlett, 1995), memories are linked using their temporal and cause-effect commonsense relations (Hwang et al., 2021). Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023). Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation, such that they provide tailored information (Chae et al., 2023) for RG (Phase III). Our contributions are two-fold:
@@
  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline augmentation for long-term conversations. Theanine outperforms representative baselines across automatic, LLM-based, and human evaluations of RG. Also, we confirm that Theanine leads to higher retrieval quality, and its procedures align with human preference. To our knowledge, we are the first to model the use of timelines (i.e., linked relevant memories) in memory management and response generation.
@@ 2 Methodologies
We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents inspired by how humans store and retrieve memories for conversations (Figure 2):
@@ 2.2 Timeline Retrieval and Timeline Refinement (Phase II)
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] We empirically set $n$ to 1, as we observe a high degree of overlap across timelines extracted from the same $C_{re}$, which can lead to redundant information (i.e., input tokens) for RG. Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$. In practice, $M_{re}$ are processed together, although processing them 1-by-1 yields the same result. Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
@@ 3 Experimental Setups
Table 1: Automatic evaluation of response quality (average of sessions).
@@ 3.1 Datasets of Long-term Conversations
There are limited datasets for long-term, multi-session conversations. Firstly, Multi-Session Chat (MSC) (Xu et al., 2022a), is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean. Recently, Jang et al. (2023) release a new dataset, Conversation Chronicles (CC). Unlike MSC, CC augments speakers with defined relationships, such as “employee and boss”. Apart from these open-domain datasets, the Psychological QA, [^8] https://www.xinli001.com/ addresses long-term conversations under clinical scenarios in Mandarin.
We opt for MSC and CC for evaluation to focus on English conversations, leaving multilingual and domain-specific conversations (e.g., DuleMon, CareCall, and Psychological QA) to future work.
@@ 3.2 Baselines
To evaluate Theanine, in addition to naive baselines that utilize all past dialogues or memories, we incorporate the following settings:
@@
Memory Update. We utilize LLMs to implement a widely used updating algorithm proposed by Bae et al. (2022) at the end of each dialogue session. This algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
RSum-LLM. An LLM-only generative method that recursively summarizes and updates the memory pool, generating responses w/o a retrieval module (Wang et al., 2023).
MemoChat. Proposed by Lu et al. (2023), it leverages LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
COMEDY. Proposed by Chen et al. (2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation. It then selects compressed memories to augment response generation.
@@ 3.3 Models and Implementation Details
Large language models. In all experiments, including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 6), (ii) memory update, and (iii) response generation. Temperature is set to $0.75$.
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity for settings involving retrievers. In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3. For the “Memory Retrieval” baseline, we set $k=6$ following Xu et al. (2022a).
Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost identical in session 1 $\sim$ 2 (no memory to update).
@@ 4 Evaluation Scheme 1: Automatic and Human Evaluations
To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation. We now present several key findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
@@
Table 1 presents the agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020). Across both datasets, Theanine, achieves superior response quality than various baselines. Although, compared to Memory Retrieval, Theanine scores slightly lower in overlap-based metrics (i.e., B-4 and R-L) in MSC, it largely outperforms Memory Retrieval in embedding-based metrics. Interestingly, including ours, methods without memory update generally yield higher scores, justifying our proposal towards an update-, removal-free memory management for lifelong dialogue agents.
To gain deeper insights into our design, we investigate the impact of removing Theanine’s relation-awareness during memory linking (Phase I-2) and Timeline Refinement (Phase II-2). Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
In Table 2, we observe a ranking in terms of contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement. This observation confirms the efficacy of constructing a memory graph with causal relations. Moreover, utilizing this graph structure to collect timelines of relevant events yields higher RG quality than conventional retrieval, despite the smaller $k$ (3 vs. 6) in initial retrieval. Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG. We leave it to future work.
@@
(Finding 4) Humans and G-Eval reveal that Theanine leads to higher retrieval quality regarding both helpfulness and accuracy.
Beyond agent responses, we further investigate how different memory construction methods affect the quality of memory retrieval. Given the same current dialogues as queries for memory retrieval, Figure 5 shows head-to-head comparisons (ours vs. baselines) regarding whose retrieved memories more effectively benefit RG. We observe higher win rates for Theanine in all comparisons, especially in human evaluations. This suggests that our method can facilitate more helpful memory augmentation for response generation.
In addition to helpfulness, objectively measuring retrieval accuracy is crucial. Since existing datasets of long-term conversations do not provide a golden mapping between dialogue contexts and memories (i.e., golden memories for retrieval), we identify 50 dialogue contexts (i.e., test instances) that require a past memory for RG, and manually measure the retrieval accuracy of different agents. The results shown in Table 3 indicate that Theanine and its ablations demonstrate higher retrieval accuracy than baselines, and the ranking here aligns with Table 1 and success rates in Table 4.
@@
Methods (Agents) Golden Memory is Retrieved/collected (%)
Memory Retrieval 68.00
+ Memory Update 64.00
MemoChat 56.00
COMEDY 48.00
[Uncaptioned image]Theanine (Ours) 72.00
Table 3: Human evaluation of the accuracy of memory retrieval (we examine 50 test instances).
@@
Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories contribute towards reliable lifelong human-agent interaction.
For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting. In Figure 6, Theanine not only leads to a small number of contradictory responses (4%) but also demonstrates the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines. We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses more directly aligned with the past. This alignment is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020). Furthermore, such entailing and non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
@@
As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, which explains its contribution to performance. Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG. Examples of Theanine’s phases and RG are in Appendix G.